[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Oct 03, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Tue Oct 3 13:29:09 UTC 2017


Here is today's summary of NHC computer operations:
--- Monica

----------

TAFB:
GOES-E images did not come in for 0515 and 0545 UTC. TAFB ran the GOES  
image retrieve program on Shiner and it successfully retrieved the 0545 UTC  
image, but not the 0515 UTC image. The 0545 UTC image was retrieved first,  
so they are not sure if that was the reason the 0515 UTC image was not  
retrieved. Matt Sardi and Monica will look into this, and provide TAFB with  
proper instructions and guidance on if there is a certain order that images  
need to be retrieved in for future reference.
HSU:
HSU had difficulty launching the GTWO from the EPAC desk, since the machine  
hung up and didn't show probabilities. Later, they attempted sending from  
tuna, which worked fine. HSU requested a hostname label on the monitor for  
the other NAWIPS system at EPAC desk for future issues. TSB will  
investigate the NAWIPS workstation and add a label as requested.
Other items:
1) Today (Oct 3rd), WCOSS machines Luna and Tide will be unavailable for a  
24-hr parallel prod test. Please let TSB know if anything strange happens  
when connecting to WCOSS as it sometimes tends to do when NCO does work on  
WCOSS.
2) Tomorrow (Oct 4th) NCO will remove the /gpfs/hps/nhc links to  
/gpfs/hps3/nhc on the WCOSS Crays. TSB has made the necessary changes to  
our code, but please let TSB know if there is any difficulty connecting to  
or sending and retrieving files to/from WCOSS this week.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171003/c271ee57/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list